charge distribution in capacitors

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
if two capitors are in series
\begin{displaymath}
C_{\rm eq} = C_1 + C_2.
\end{displaymath}
it two capacitors in parallel
\begin{displaymath}
\frac{1}{C_{\rm eq}} = \frac{1}{C_1} + \frac{1}{C_2}.
\end{displaymath}


sher mohammad
askiitians faculty
b.tech, iit delhi